ID |
Date |
Author |
Type |
Category |
Subject |
9018
|
Fri Aug 16 13:25:50 2013 |
Koji | Update | ASS | ASX model/screen cleaning up |
[Koji Manasa]
Yesterday we cleaned up the ASX model and screens to have more straight forward structure of the screen
and the channel names, and to correct mistakes in the model/screens.
The true motivation is that I suspect the excess LF noise of the X arm ALS can be caused by misalignment
and beam jitter coupling to the intensity noise of the beat. I wanted to see how the noise is affected by the alignment.
Currently X-end green is highly misaligned in pitch.
- Any string "XEND" was replaced by "XARM", as many components in the system is not localized at the end table.
- The name like "XARM-ITMX" was changed to "XARM-ITM". This makes easier to create the corresponding model for the other arm.
- There was some inconsistency between the MEDM screens and the ASX model. This was fixed.
- A template StripTool screen was created. It is currently saved in users/koji/template as ASX.stp.
It will be moved to the script directory once it's usefulness is confirmed.
The next step is to go to the end table and manually adjust M2 mirror while M1 is controlled by the ASX.
The test mass dithering provides the error signal for this adjustment but the range of the PZT is not enough
to make the input spot position to be controlled. In the end, we need different kind of matching optics
in order to control the spot position. (But is that what we want? That makes any PZT drift significantly moves the beam.) |
9017
|
Fri Aug 16 09:35:18 2013 |
Steve | Update | VAC | Vacuum Normal state recognition is back |
Quote: |
Quote: |
Quote: |
Quote: |
Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr. All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open. One of the roughing pumps was also turned on. Other stuff we didn't notice? Bad.
|
Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind. All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.
|
CLOSE CALL on the vacuum system:
Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.
The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly accumulated to reach ~10-50 Torr
I disconnected_ immobilized and labelled the following 6 valves: the 4 large ion pump gate valves and VC1, VC2 of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!
This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.
The down side of this that the computer can not identify vacuum states any longer.
This vacuum system badly needs an upgrade. I will make a list.
|
While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.
All valve closed to default condition. I realized that the current indicator was red at 0.2A and the voltage fluctuated from 3-13V
Increased current limiter to 0.4A and set voltage to 24V I think this was the reason for the caos of valve switching during the VME swap.
|
Based on the facts above I reconnected VC1 and VC2 valves. State recognition is working. Ion pumps are turned off and their gate valves are disabled.
We learned that even with closed off gate valves while at atmosphere ion pumps outgass hydrocarbons at 1e-6 Torr level. We have not used them for this reason in the passed 9 rears.
I need help with implementing V1 interlock triggered by Maglev failure signal and-or P2 pressure.
MEDM screen agrees with vacuum rack signs. |
Attachment 1: VacuumNormal.png
|
|
Attachment 2: vacValvesDisabled.jpg
|
|
9016
|
Thu Aug 15 21:42:53 2013 |
Charles | Update | ISS | ISS - Schematic + PCB Layout |
After many, many moons of getting to know exactly how frustrating Altium can be, I have completed the PCB layout for my ISS board (final page of ISS_v3.pdf).
Before I get into detail about the PCB, there is one significant schematic change to note: the comparator circuit was changed (with significant help from Koji) so that the voltage reference for boost triggering is established in a more logical way. Instead of the somewhat convoluted topology I had before, now there are only two feedback resistors, R82 and R83. Because their resistances (500k and 50k respectively) are so much larger than the total resistance of the 1k potentiometer (used to establish a tunable threshold voltage), the current flowing through the feedback loop is negligible compared to the 5 mA current flowing through the potentiometer (the pot is rated for 2 W and with 5 mA -> 25 mW dissapation). This allows one to set the threshold voltage for my schmitt trigger, at pin 2 of both the pot and the comparator, entirely with the pot. This trigger also has hysteresis given by the relation deltaV ~ (R83/R82) * (Voh - Vol) where deltaV is the separation between threshold voltages, Voh is the high-level comparator ouput and Vol is the low-level comparator output. Koji simulated this using CircuitLab and I plan to verify the behavior by making a quick prototype circuit.
Now, on to the PCB. The board itself is of a 'standard' LIGO size (11" x 6") has 3 routing layers and 3 internal planes, one for +15 V, one for -15 V and one for GND. In the attached pdf, red is the top routing layer, blue is the bottom layer and brown is the middle routing layer (used for ±5 V exclusively). The grey circles are pads and vias (drilled through) and anything in black is silkscreen overlay. I placed each component and track by hand, attempting to minimize the signal path and following the general rules below,
- Headers for power, ±5 V and ±15V, are at the back of the board
- For sections of the board such as filter stages or buffers, resistors and capacitors were grouped around their respective op-amps.
- As often as was possible, routing was confined to the top layer. Tracks on the bottom layer were placed mostly out of necessity (i.e. no possible connection on top routing layer).
- The signal generally proceeds from left to right (directions with respect to the attached printout) in the same logical order as on the schematic sheets. Refer to the global sheet (page 1) of the attached "ISS_v3.pdf".
- External ports such as the PD input, various monitoring ports and panel mounted switches/LEDs were all connected to the board via headers located along the front edge. These are also ordered following the schematic layout.
- Occasionally, similar signal paths were grouped together although this was a rarity on my board
Sections of the board have been partitioned and labeled with silkscreen overlay to help in both signal pathway recognition as well as eventual troubleshooting.
On the board, I have also included holes so that it can be mounted inside of an enclosure. There is a DCC number printed as well as a 'barcode' (TrueType font: IDAutomationC39S), although they both contain filler asterisks as I haven't published this to the DCC and thus do not have a number. |
Attachment 1: ISS_v3.pdf
|
|
Attachment 2: ISS_v3-Power_Reg.pdf
|
|
9015
|
Thu Aug 15 19:05:07 2013 |
manasa | Update | Green Locking | ALS out of loop noise |
Beat notes were recovered for both the arms.
I locked the arms to IR using PDH and measured the ALS out of loop noise at the phase tracker output.
The Y arm has the same 300Hz/rtHz rms. The X arm rms noise measures nearly the same as the Y arm in the 5-500Hz region (X arm has improved nearly 10 times after the last whitening filter stage change old elog ).
The noise in the ALSX error signals could be related to the bad alignment and conditions at the X end. |
Attachment 1: ALS_OutLoop.pdf
|
|
9014
|
Thu Aug 15 12:30:17 2013 |
manasa | Update | Green Locking | Lost beat notes |
[Koji, Nic, Manasa]
Update from last night.
Koji and I realigned the green optics on the PSL to start working on the ALS.
We set on a beat note search. We couldn't find the beat note between any of the arm green transmission and the PSL green. All we could see was the beat between the X arm and the Y arm green leakage.
Since we had the beatnote between the 2 green transmission beams, we decided to scan the PSl temperature. We scanned the SLOW actuator adjust of PSL; but couldn't locate any beat note. The search will continue again today. |
9013
|
Thu Aug 15 09:34:12 2013 |
Steve | Update | General | Wilcoxon cables rescued |
Eric and Steve,
We removed Wilcoxon Accelerometer PS and Amplifier unit under the BS optical tabel yesterday. The six cabels going to DAQ were labeled and left in place. Gain setting were 100, except channel 3 was 10.
The ~ 40 m long 2 sets of 3 cables were very happy to get their kinks out. Especially the set going just south of ITMX optical table.
We have to take better care of these cables! Your data will be useless this way. |
Attachment 1: rescuedGraycables.jpg
|
|
Attachment 2: wilconoxOut.jpg
|
|
Attachment 3: chanGains.jpg
|
|
9012
|
Thu Aug 15 01:51:50 2013 |
Koji | Summary | General | RFM<->Dolphin bridge distributed to c1rfm and c1mcs |
Since the RFM-Dolphin bridges for the ASX model was added to the c1rfm model, c1rfm kept timing-out from the single sample time of 60us.
The model had 19 dolphin accesses, 21 RFM accesses, and 9 shared memory (SHM) accesses.
At the beginning 2 RFM and 2 SHM accesses were moved to c1sus (i.e. they were mistakenly placed on c1rfm).
But this actually made the c1sus model timed out. So the model was reverted.
The current configuration is that the WFS related bridges were accommdated in the c1mcs model.
This made the timing of c1rfm ~40us. So it is safe now.
On the other hand, the c1mcs model has the time consumption of ~59us. This is marginal now.
We need to understand why any RFM access takes such huge delay. |
9011
|
Wed Aug 14 08:24:20 2013 |
Steve | Update | VAC | pumpdown at day 8 |
|
Attachment 1: pd76md8.png
|
|
9010
|
Tue Aug 13 22:21:12 2013 |
Koji | Summary | General | Minicircuit Filter TFs (AG4395A test) |
As a part of the network analyzer test in the previous entry, the transfer functions of Mini-Circuits filters we have at the 40m were measured.
<<List of the filters>>
- LPF (SMA): SLP1.9, SLP5, SLP21.4, SLP30, SLP50, SLP100, SLP150, SLP750
- LPF (BNC): BLP1.9, BLP2_5, BLP5, BLP30
- BPF (SMA): SBP10.7, SBP21.4, SBP70
- HPF (SMA): SHP25, SHP100, SHP150, SHP200, SHP500
|
Attachment 1: Minicircuit_LPF.pdf
|
|
Attachment 2: Minicircuit_BPF.pdf
|
|
Attachment 3: Minicircuit_HPF.pdf
|
|
Attachment 4: 130813.zip
|
9009
|
Tue Aug 13 21:49:32 2013 |
Koji | Summary | General | Testing new AG4395A network analyzer |
New AG4395, sn MY41101114 for West Bridge Labs was delivered. For the test purpose it is at the 40m now.
I made a series of tests in order to find anything broken.
Network analyzer test
- RF out / Rch test
RF out directly connected to R input channel.
The received power at the R-ch was measured while the output was swept from 10Hz to 500MHz.
The RF power was changed from -50dBm to +15dBm with +10dBm increment (but the last one).
The attenuator setting was changed from 50dB to 0dB.
=> The configured output power was properly detected by the R channel.
=> RF output is producing the signal properly. R-ch is detecting the produced signal properly.
- Ach/Bch test
Same test as above for Ach and Bch
=> Same result as above
=> A-ch and B-ch are detecting the produced signal properly.
- Transfer function test
Connect a power splitter to the RF out. Detect the split signals by R-ch and A-ch
=> Measurement is at around 0dB +/- 1dB up to 500MHz.
Same measurement for B-ch
=> Same result
=> A/R and B/R indicates proper transfer function measurements.
- Calibration
RF out was split in to two. One was connected to R-ch. The other was connected to A-ch.
The thru response calibration was run.
=> The thru calibration was performed properly.
- Practical tranfer function measurements.
In the above calibration setup, various RF filters were inserted in the Ach path.
The measured data was extracted via GPIB connection.
=> Practical transfer function measurements were performed.
=> GPIB connectivity was confirmed
External reference test
- External 10MHz reference from an SRS frequency counter was connected to Ext Ref In
=> Ext Ref indicator on the screen appeard
=> The internal oscillator seemed to be locked to the external reference in
Spectrum analyzer test
- Measured the signals from DS345 by R/A/B ch
Sinusoidal signal (1V) swept from 10MHz to 30Mhz
=> Corresponding moving peak was detected in each case
- Noise level measurement
R/A/B channels were terminated. The attenuation at each port was set to 0dB.
Frequency span was changed between 500MHz, 10MHz, 100kHz, 1kHz.
=> Noise level of ~10nV/rtHz between 0.1-500MHz was confirmed. All R/A/B channels have the same performance.
|
Attachment 1: AG4395A_noise.pdf
|
|
9008
|
Tue Aug 13 21:09:03 2013 |
manasa | Update | Green Locking | Arms ready for ALS |
I aligned both the X and Y end green to the arms.
The transmitted green were aligned at the PSL table green optics to the beat PDs.
Beat notes were retrieved.
To do:
1. Check Y arm ALS with previous performance.
2. Troubleshoot X arm ALS.
3. Edit the automation scripts for ALS.
4. Modify ALS model to talk to LSC instead of suspension models. |
9007
|
Tue Aug 13 17:20:54 2013 |
Koji | Update | CDS | [Fixed] c1iscex needs help |
c1x01 timing issue was solved. Now all of the models on c1iscex are nicely running.
Symptons
- c1x01 was synchronized to 1PPS in stead of TDS
- C1:DAQ-DC0_C1X01_STATUS (Upper right indicator) was red. The bits were 0x4000 or 0x2bad.
C1:DAQ-DC0_C1X01_CRC_SUM kept increasing
- c1scx, c1spx, c1asx could not get started.
Solution
- login to c1iscex "ssh c1iscex "
- Run "sudo shutdown -h now "
- Walk down to the x end rack
- Make sure the supply voltages for the electronics are correct (See Steve's entry)
- Make sure the machine is already shutdown.
- Unplug two AC power supply of the machine.
- Turn off the front panel switch of the IO chassis
- Wait for 10sec
- Turn on the IO chassis
- Plug the AC power supply cables to the machine
- Push the power switch of the realtime machine |
9006
|
Tue Aug 13 13:30:41 2013 |
Alex Cole | Configuration | Electronics | Cable Routing |
I routed cables (RG405 SMA-SMA) from several demodulator boards in rack 1Y2 to the RF Switch in rack 1Y1 using the overhead track. Our switch chassis contains two 8x1 switches. The COM of the "right" switch goes to channel 7 of the "left" switch to effectively form a 16x1 switch. The following is a table of correspondences between PD and RF Switch input.
PD |
Left/Right Switch |
Channel Number |
REFL11 |
R
|
0 |
POX11 |
L |
0 |
AS55 |
R |
1 |
REFL55 |
R |
7 |
POP22 |
R |
6 |
REFL165 |
R |
5 |
REFL33 |
L |
7 |
ThePOP110 demod board has not yet had a cable routed from it to the switch because I ran out of RG405.
We should also consider how important it is to include MCREFL in our setup. Doing so would require fabrication of a ~70 ft RG405 cable. |
Attachment 1: photo_(6).JPG
|
|
9005
|
Tue Aug 13 11:54:40 2013 |
Alex Cole | HowTo | Electronics | RF PD Fiber-Coupled Laser Operation |
This post pertains to the fiber-coupled diode laser mounted in rack 1Y1.
To turn the laser on, first turn the power supply's key (red) to the clockwise. Then make sure that the laser is in "current" mode by checking that the LED next to "I" in the "Laser Mode" box in lit up. If the light is not on, press the button to the right of the "I" light until it is. Now press the output button (green). This is like removing the safety for the laser. Then turn the dial (blue) until you have your desired current. Presently, the current limit is set to around 92 mA.
To turn the laser off, dial the current back down to 0mA and turn the key (red) counterclockwise. |
Attachment 1: photo_(4).pdf
|
|
9004
|
Tue Aug 13 11:40:19 2013 |
Alex Cole | Summary | Electronics | RFPD Demod Filter Frequency Response Measurement |
For the RF PD Frequency Response Measurement project, we get each PD signal from the "PD RF Mon" output of each demodulator board corresponding to our PD under test. Therefore we can't neglect the frequency response of various filters inside the demodulator board. I used our Agilent 4395 Network Analyzer to gather frequency response data for each demodulator board being considered for the RFPD frequency response project (AS55, REFL11, REFL33, REFL55, REFL165, POX11, POP22, POP110).
The NA swept over a frequency range of 1-500 MHz. Data was collected using NWAG4395A (from the netgpibdata directory). It should be noted that the command line options -a 16 -x 15 (averaging=16 and excitation amplitude=15 dBm[the max]), in addition to the usual command line options described in the help file, were used to minimize noise.
The data is located in /users/alex.cole. The file names are in the format [PDNAME]DemodFilt_1000000.dat (e.g. REFL11DemodFilt_1000000.dat). Results for POP110 are shown below. |
Attachment 1: photo_(3).JPG
|
|
Attachment 2: test.jpg
|
|
9003
|
Tue Aug 13 11:04:44 2013 |
Steve | Update | PEM | fluorecent lights |
Our fluorecent lights became obsolete. We'll have change fixtures over to some more energy efficient one. Do you have any recommendation regarding to less noise performer unit?
We may go this direction of LED fluorecent lamps ? |
9002
|
Tue Aug 13 07:40:53 2013 |
Steve | Update | CDS | c1iscex needs help |
Sorrensen ps ouput of +15V at rack 1X9 was current limited to 10.3V @ 2A
Increased threshold to 2.1A and the voltage is up to 14.7V |
Attachment 1: c1iscexSick.png
|
|
9001
|
Mon Aug 12 23:13:14 2013 |
Jenne | Update | ASC | PRCL ASS software in place |
I guess I was thinking that POPDC was a proxy for any type of PRCL lock. Even if we're sideband locked, there is still some signal in POPDC (although it is very small relative to a carrier lock - ~40cts vs. 1,000cts). However, as soon as this question was asked of me, I realized that one of the 2f demodulated signals made more sense.
Since I want the ability to choose between POP110 and POP22, I have put a little 1x3 input matrix before the PRCL lockins in the ASS model. Since POPDC was already there, I included it as an option in the matrix (in case we ever want to do some PRCL ASS after we have some carrier resonating as well). |
9000
|
Mon Aug 12 21:27:03 2013 |
manasa | Update | CDS | c1iscex needs help |
I started to modify the c1asx model to reduce the RFM model from hitting its max time.
Instead of bringing in ASS, I have modified ASX to do everything and only the clock signals to ITMX pitch and yaw are now going through RFM. RFM is still hitting 62usec and I suppose that is because of the problems with c1iscex.
c1iscex not happy
Cause and symptoms
While restarting the models, c1iscex crashed a couple of times because of some errors and had to be powercycled. The models were modified and they seem to start ok.
But it looks like there is something wrong with c1iscex since the models were started. The GPS time is off and C1:DAQ-DC0_C1X01_CRC_SUM keeps building up even for c1x01 which was left untouched.
Trial treatments
1. Since c1x01 ans c1spx were not touched,c1scx and c1asx were killed and we tried to start the other models. This did not help.
2. Koji did a manual daqd restart which did not help either.
We are leaving c1iscex as is for the time being and calling Jamie for help.
P.S. While making the models, I had created IPCx_PCIE blocks in c1iscex which do not exist. I changed them to RFM and SHMEM blocks. This did not allow me to compile the model and was only spitting errors of IPCx mismatch. After some struggle and elog search I figured out from an old elog that eventhough the IPCx blocks are changed in the model, the old junk exists in the ipc file in chans directory. I deleted all junk channels related to the ASX model. The model compiled right away. |
8999
|
Mon Aug 12 17:30:03 2013 |
Koji | Update | ASC | PRCL ASS software in place |
Why POPDC??? |
8998
|
Mon Aug 12 15:39:40 2013 |
Jenne | Update | ASC | PRCL ASS screens in place |
I have added the PRCL ASS to the main ASS screen, and created the servo and lockin screens. The filters loaded are the same as those used for the arms (bandpasses and lowpasses for the lockins, and an integrator for the servo).
I'm going to try to lock, and get the ASS to work. |
8997
|
Mon Aug 12 14:05:34 2013 |
Jenne | Update | ASC | PRCL ASS software in place |
Quote: |
- We are ready to implement ASS for PRM
|
I have added an IPC sender from the LSC model, to send POPDC to ASS. I have copied over the structure of the arms' ASS, to do the same for PRCL. I have set it up to dither the PRM, and feed back to the PRM. I did not include an LSC set, since I'm assuming that we'll set the input pointing with the arms, and just want to move the PRM to maximize POPDC.
Models have been compiled, installed, and restarted, and the daqd was restarted. |
8996
|
Mon Aug 12 13:30:33 2013 |
Jamie | Update | CDS | X-End Green ASS - Roundup |
Quote: |
I'm not really sure why the ASS was involved in this. I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner. Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model. I had thought that this was expressly forbidden (although I'm not sure why). I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level.
|
I'm not sure if it's forbidden by the RCG, but you should definitely NOT do it. All IO, whether it be between ADC/DACs or IPCs, should always be at the model top level. That's what keeps things portable, and makes it easier to keep track of where are signals are going/coming from. |
8995
|
Mon Aug 12 12:57:59 2013 |
Jenne | Update | CDS | X-End Green ASS - Roundup |
Quote: |
- The SIMULINK model has been modified to accommodate an option to dither the cavity mirrors and not the PZT mirrors. Details are as follows:
- I have sent the LO signals (CLK,SIN and COS) from the ASS model to the ASX model via the RFM model. Appropriate changes were made to all these three models, and recompiling and restarting the models was done without issue. The SIN and COS signals are used to demodulate green transmission at the dither frequencies. ***The CLK signal is not required to be sent between models as it is not being used by ASX (I turn the dither ON using the channels already set up for ASS). I realised this a little late, and at present the ASS and RFM models are compiled such that the CLK signal is also sent from ASS to RFM. This can be removed, thus freeing up 4 unnecessary inter-process communication channels. Also, I am not too sure if this is relevant, but the maximum computation time of both the RFM and ASX models seem to have gone up after I added these inter-process communication links.***
|
Getting rid of the LO transmission will certainly help / be good. After adding these channels, the RFM model is regularly hitting 62usec (out of a max acceptable of 60).
I'm not really sure why the ASS was involved in this. I feel like it might have been simpler to just do everything in the ASX model, to keep things cleaner. Also, the IPC blocks for this stuff (in both ASS and ASX) are not on the top level of the model. I had thought that this was expressly forbidden (although I'm not sure why). I'm emailing Jamie, to see if he remembers what, if anything, is breakable if the IPC blocks are down a level. |
8994
|
Mon Aug 12 10:44:22 2013 |
rana | Update | LSC | PRMI(sb) lock recovered |
In the past, we used to use Stefan's 'ezcademod' or Matt's 'ezlockin' to do auto phase adjustment.
JoeB / Jamie are working on python replacements for these tools, but in the near term possibly I can make a bash script to use ezcaservo and the existing LOCKINs to do this. |
8993
|
Sat Aug 10 05:53:51 2013 |
gautam | Update | CDS | X-End Green ASS - Roundup |
Over the last three days, I've had the interferometer to test and optimize the ASX Servo. Based on what I have seen, I think the conclusion is that with the current parameters, the servo does its job provided the input pointing set up at the endtable with the coarse adjustment knobs is reasonably good. Once the cavity is aligned and IR transmission maximized using ASS, I have been able to get the green transmission up to 0.8 which is close to the best we had pre-vent. I have not been elogging regularly over the last few days, so this one is going to be a longish one.
Major changes made:
- The SIMULINK model has been modified to accommodate an option to dither the cavity mirrors and not the PZT mirrors. Details are as follows:
- I have sent the LO signals (CLK,SIN and COS) from the ASS model to the ASX model via the RFM model. Appropriate changes were made to all these three models, and recompiling and restarting the models was done without issue. The SIN and COS signals are used to demodulate green transmission at the dither frequencies. ***The CLK signal is not required to be sent between models as it is not being used by ASX (I turn the dither ON using the channels already set up for ASS). I realised this a little late, and at present the ASS and RFM models are compiled such that the CLK signal is also sent from ASS to RFM. This can be removed, thus freeing up 4 unnecessary inter-process communication channels. Also, I am not too sure if this is relevant, but the maximum computation time of both the RFM and ASX models seem to have gone up after I added these inter-process communication links.***
- The rest of this part of the servo is a replica of the part where PZT mirrors are dithered. At present the servo output is the sum of its two branches (PZT mirror dither branch and cavity mirror dither branch) which works fine under the assumption that at any one time, only one arm will run. Ideally, the summing block should be replaced by a switch. However, when I tried (in an earlier attempt to include the cavity dither) to do this and restart the model, c1iscex crashed, and so I decided against using the switch block for this trial.
- The control signal generated using green transmission demodulated at the ETM dither frequencies are used to actuate on M1 while the ITM ones are used to actuate on M2. Of course, by setting the appropriate off-diagonal elements in the output matrix, this can be modified as desired.
- The main MEDM screen has been updated to reflect the new additions to the SIMULINK model. Screenshot is attached. The picture isn't entirely accurate as the monitor channels in the upper row actually show the servo output + slider output. This needs to be changed in the model, and a new set of monitors need to be added to the MEDM screen. In the end, we require four sets of monitor-points in the model: PZT dither servo output, cavity dither servo output, sum of these with any offset from the PZT sliders, and the sum of the latter with the dither signal (this is what eventually goes to the PZT mirrors while the dither is on).
- I added scripts to the MEDM screen that turn the PZT mirror dither servo on and off. Note that when you want to run a new script on an MEDM screen using medmrun, you need to change the permissions of the file by going to the path where your script is located and running chmod 755 <name of script>. Manasa has updated the same on the wiki.
Details of tests runs:
For the most part, I have been trying to optimize the PZT mirror dither servo. To this end, I did the following:
- Went to the X-end and fixed the input pointing which was not optimal. Manasa first aligned the arm and ran ASS to maximize the IR transmission. I then used the coarse adjustment knobs on the mirror mounts to get the green transmission up to ~0.6.
- I then set the following parameters in the servo (these are all in the script, path to which is /opt/rtcds/caltech/c1/scripts/ASX):
- LO frequencies of 10, 19, 34 and 39 Hz respectively for M1 PIT, M1 YAW, M2 PIT and M2 YAW.
- LO amplitudes of 75 for all the four degrees of freedom (determined by using PZT calibration to see what amplitude would couple 10% of power into the first higher-order-mode assuming a perfectly aligned beam to start with.
- LIA BP filters centered at the above frequencies with 2Hz passband on either side.
- LIA LP filters with corner frequency 0.5 Hz.
- LIA Signal filter bank gain set to 100 for all degrees of freedom.
- LIA Demod I phase filter bank gain set to 5 for all degrees of freedom.
- Control filter gains to 1 for all degrees of freedom (control filters are all integrators).
- Demod phase set to 0 for all degrees of freedom. I did not really try to optimize this but the servo seems to be doing reasonably well even with this setting.
- Overall servo gain to 1 (the servo worked well when I increased this to 5 as well, but became unstable when I increased it further).
- I ran the servo. Observations were as follows:
- Having fixed the input pointing to get green transmission up to ~0.6, the servo was able to improve it to ~0.8, which is the best we had after hours spent at the X-end prior to the vent.
- Given a good input pointing, we can use the PZT mirrors to lock to 00 mode from some misaligned state using either the sliders, or by leaving the servo on, and helping it out at the points where it gets 'stuck' in some higher mode using either the sliders or by toggling the shutter.
- In order to recover green transmission of ~0.8, it was often necessary to first run ASS to optimize the IR transmission. Otherwise, green-transmission saturates at ~0.6 or 0.4 depending on the misalignment of the arm cavity mirrors. The servo was unable to change the input pointing enough to deal with overly misaligned cavity mirrors.
- The servo is sometimes capable of bringing about mode-hopping from a higher order mode to a lower one, though this is not always the case as the PDH lock is sometimes too strong, in which case I toggled the shutter after which the servo kicked in.
- I tested the servo under as many different conditions as I could. For instance, having left the green shutter open overnight, I saw that the transmission had fallen from 0.8 (which was what we saw on Thursday night) to ~0.4 on Friday morning. Running the servo got the transmission up to 0.6. I then asked Manasa to run ASS, (while leaving the ASX servo on), after which point the green transmission went up to 0.8. Sometimes, the servo locks to a 'bad' 00 mdoe, where the transmission saturates at ~0.2, but toggling the shutter fixes this most of the time.
Attempt to measure transfer function:
One of the things that came up during my presentation was how fast the loop was capable of responding. I was able to get a quantitative idea of this by playing around with the overall servo gain. Initially, it took ~30 seconds for the servo to get the transmission up to its peak value, with a servo gain of 1. When I ramped this up to 5, the response was much faster, with the peak transmission being reached in ~5seconds.
I wanted to get a more quantitative picture, and hence tried to measure the transfer function by first injecting an excitation into the 'SIG' filter-bank in the demodulation stage. However, coherence between the IN1 and IN2 signals was very poor for all the amplitude configurations I tried. At Jenne's suggestion, I tried injecting the excitation at the control-filters stage, but found no improvement. Perhaps the amplitude envelope was wrong or the measurement technique has to be rethought.
Misc remarks:
- M1 is the first steering mirror and M2 is the second one (right before the beam enters the arm cavity).
- Though I have added the cavity dither feature to the model, I was not able to optimize this servo. Some calculations need to be done to get an estimate of the output matrix, after which the filter gains etc can be optimized.
- Today, I cleaned up my temporary setup at the SP table to calibrate the PZTs. Most of the hardware for the Y-end is now in the tupperware box. The QPD and laser have been restored to the optical bench next to MC2 where I found them. The second KEPCO HV supply which I had set up has now been installed at 1Y4 in anticipation of the PZT mirrors at the Y-endtables. It is currently powered OFF.
- Performance plots to follow as I have not pulled the data out yet.
- I had bought a cake from chandler today in an effort to clear my meal plan, but in the rush in the afternoon, completely forgot about it. It is in the fridge, and is strawberry tart, hope it tastes good.
New MEDM screen:
|
8992
|
Fri Aug 9 22:51:37 2013 |
Koji | Update | LSC | PRMI(sb) lock recovered |
PRMI(sb) lock was recovered
PRMI lock
- Stared at the time series data of the REFL demod signals, and decided to use REFL165I&Q for the locking.
- Jiggled the demodulation phase of REFL165 and POP110. Changed the servo gains.
- Finally found a short lock. Further optimized the parameters.
- PRM ASC was turned on by giving the identity matrices for the input and output matrices.
Now just hitting the up button is sufficient to engage the ASC servo.
- Under the presence of the ASC, the PRMI is indefinitely locked as before.
- Reacquisition is also instantaneous. (It acquires even if the ASC is left "on".)
- Actually the lock is somewhat robust even when the PRM ASC is not used.
This is VERY GOOD as we can skip one of the steps necessary for the full lock.
Although, the seismic on Friday night is very quiet.
The spot motion at POP seems to be somewhat pitch/yaw mixed, in stead of previous "totally-dominated-by-yaw" situation.
- We are ready to implement ASS for PRM
Demod phase adjustment
- Shook PRM at 580Hz / 100cnt
- Swept the demod phase of REFL165 such that the PRM peak is minimized in the Q signal
- Open DTT. Measured transfer functions between REFL165I and the Q signals of each PD.
- Minimized the PRCL signal coupling in the signals.
- The resolution of the adjustment was ~1deg.
Locking test with PRM/BS
Tried the lock acquisition only with PRM and BS. (cf. http://nodus.ligo.caltech.edu:8080/40m/8816)
This just worked nicely.
Today's locking parameters:
PRMI(sb) lock:
MC Trans: 17500
POP110I (in lock): 150
PRCL Source: REFL165(I) 106deg / 45dB / Normalization SQRT(10 POP110I) / Input MTRX 1.0
PRCL Trigger: POP110I x 1.0 50up 25down
PRCL Servo: G=+3.5 Acq: FM4/FM5 Opr: FM2/FM3/FM6/FM7
PRCL Actuator: PRM +1.0
MICH Source: REFL165(Q) 106deg / 45dB / Normalization SQRT(0.1 POP110I) / Input MTRX 1.0
MICH Trigger: POP110I x 1.0 50up 25down
MICH Servo: G=-10 Acq: FM4/FM5 Opr: FM2/FM3/FM6
MICH Actuator: (ITMX -1.0 / ITMY +1.0) or (BS 0.5 / PRM -0.267)
Demod phases:
AS55 -17deg
REFL11 135deg
REFL33 -18deg
REFL55 120deg
REFL165 106deg |
8991
|
Fri Aug 9 21:05:28 2013 |
Koji | Update | SUS | fixed: SRM coils fine - problem with slow bias slider |
Now the SRM Yaw bias in yaw is functional without any strage behavior.
The problem was found at the connector of the flat ribbon cable from the DAC to the cross connect.
I used the extender board to diagnose the SRM coil driver circuit at 1X4.
The UL coil input did not show any sign of voltage no matter how the bias slider was jiggled.
I opened the side panel of the rack and found the signal was absent at the cross connect which relays two flat ribbon cables
for the SRM coil driver. I checked the DAC output with a multimeter. All the bias outputs were OK at the DAC.
Then I opened the IDC connector at the DAC side of the crossconnect as the signal was already missing there.
I found that the flat ribbon cable was a half line shifted from the supposed location.
This resulted a short circuit of the DAQ +/- pins for the SRM UL coil.
I recrimped the connector and now the SRM Yaw slider is back.
This changed the nominal position of the SRM. The new slider values were saved. |
8990
|
Fri Aug 9 16:49:35 2013 |
Jenne, manasa | Update | Electronics | Post-vent alignment cont'd - RFPDs |
Notes to the fiber team:
I am aligning beam onto the RFPDs (I have finished all 4 REFL diodes, and AS55), in preparation for locking.
In doing so, I have noticed that the fiber lasers for the RFPD testing are always illuminating the photodiodes! This seems bad! Ack!
For now, I blocked the laser light coming from the fiber, did my alignment, then removed my blocks. The exception is REFL55, which I have left an aluminum beam dump, so that we can use REFL55 for PRM-ITMY locking, so I can align the POP diodes.
EDIT: I have also aligned POP QPD, and POP110/22. The fiber launcher for POP110 was not tight in its mount, so when I went to put a beam block in front of it and touched the mount, the whole thing spun a little bit. Now the fiber to POP110 is totally misaligned, and should be realigned.
What was done for the alignment:
1. Aligned the arms (ran ASS).
2. Aligned the beam to all the REFL and AS PDs.
3. Misaligned the ETMs and ITMX.
4. Locked PRM+ITMY using REFL11.
The following were modified to enable locking
(1) PRCL gain changed from +2.0 to -12.
(2) Power normalization matrix for PRCL changed from +10.0 to 0.
(3) FM3 in PRCL servo filter module was turned OFF.
5. POP PDs were aligned. |
8989
|
Thu Aug 8 21:25:36 2013 |
Koji | Update | General | Post-vent alignment cont'd |
- IPANG aligned on the QPD. The beam seems to be partially clipped in the chamber.
- Oplev of the IFO mirrors are aligned.
- After the oplev alignment, ITMX Yaw oplev servo started to oscillate. Reduced the gain from -50 to -20. |
8988
|
Thu Aug 8 18:47:41 2013 |
Sujan | Summary | PEM | Using weiner filters for subtracting signals MC_L and GUR2_X |
I used MC_L signal from the Mode Cleaner as the desired signal with GUR2_X as witness signals. I observed good subtraction where coherence is high. But there was noise added in other frequency bands. I am not sure how to avoid that.
Please find attached documents that contains relevant plots. |
Attachment 1: Results.zip
|
8987
|
Thu Aug 8 18:43:12 2013 |
Sujan | Summary | PEM | Optimally subtracting signals from two seismometers. |
An exercise of optimally subtracting one seismometer signal by another using weiner filters was done. Results have been summarized document attached. |
Attachment 1: Results.pdf
|
|
8986
|
Thu Aug 8 11:18:46 2013 |
manasa | Update | Computer Scripts / Programs | Unused scripts in ASS moved |
I was receiving missing path error when I was trying to measure the MC spot positions. Jenne pointed out that Koji had moved all the unused scripts in scripts/ASS to /scripts/ASS/OBSOLETE yesterday and in the process one of the scripts that the MC spot position measurement script calls for (MeasureSpotPositions.py) must have also been moved to the OBSOLETE directory. I moved the script to /scripts/ASS/MC so that we know the script is being used and also changed its path in the main script. |
8985
|
Thu Aug 8 10:31:28 2013 |
Steve | Update | VAC | few reminders of this vent |
1, Vacuum envelope grounds must be connected all times! After door removal reconnect both cables immediately.
2, The crane folding had a new issue of getting cut as picture shows.
3, Too much oplev light is scattered. This picture was taken just before we put on the heavy door.
4, We were unprepared to hold the smaller side chamber door 29" od of the IOC
5, Silicon bronze 1/2-13 nuts for chamber doors will be replaced. They are not smooth turning.
|
Attachment 1: GROUND!.jpg
|
|
Attachment 2: bad_Folding.jpg
|
|
Attachment 3: toomuchred.jpg
|
|
Attachment 4: BETTERholder.jpg
|
|
8984
|
Thu Aug 8 07:57:23 2013 |
Steve | Update | VAC | pumpdown at 48 hrs |
Quote: |
TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.
|
|
Attachment 1: 48hrsPd76.png
|
|
8983
|
Wed Aug 7 23:40:49 2013 |
gautam | Update | CDS | X-End Green ASS - A first update |
I have done some preliminary testing of the X-End Green ASS Servo. I will write a more detailed elog about this soon, but I thought I'd note down the important stuff here.
Yesterday, while we were venting, I aligned the X-arm to the green using the sliders on IFOalign, maximizing the transmission. Then I retook a power spectrum so as to determine the LO frequencies. Jenne pointed out that LO frequencies should not be integers (it usually suffices to append a .098725 or something to the frequency) so I made the necessary changes.
I did a first run of the servo yesterday, and more runs today. Notable points:
- I was able to lock to 00 from a 08 or 09 mode using the PZT sliders
- The green transmission having locked to 00 was ~0.2. I then ran the servo and got it up to ~0.4 and then 0.6 (see time series plot attached). The servo was able to recover this level of transmission after misaligning the steering mirrors using the PZT sliders.
- This was not the optimal transmission level as when Koji moved ETMX a little, the transmission improved.
- The actuators are degenerate. Most of the time, only two of the four servos are doing anything significant. This is probably because of the fact that the two steering mirrors are so close to each other, that moving one or the other produces virtually the same effect. I do however have some cool videos of mode-hopping :)
- The range of actuation of the PZTs is probably not enough to maximize the green transmission from an arbitrary state because of point 4 (i.e. we need to move one mirror in some direction a lot, and move the other a lot to compensate for the change, and the overall gain in input pointing/alignment is marginal). It may be that things will be slightly better at the Y-end. It would also be interesting to see if there is any improvement in the servo performance by dithering the cavity mirrors as opposed to the PZT mirrors.
- To this end, I tried modifying the c1asx model to incorporate an option to dither the cavity mirrors. The plan was to make a second set of LOs in the model that output to ITMX and ETMX suspensions. However, for some reason, when I recompiled the model and restarted it, c1iscex crashed. Parity has now been restored. Note that in order to accommodate the new LOs, I had to make changes to C1SUS, C1RFM and C1SCX as well. I have since removed all my additions, saved, built and installed these models, but have not restarted them (with the exception of C1SCX which restarted when I manually restarted c1iscex).
- The plan tomorrow is to try incorporating cavity dither into the model again. This time, I'll try grabbing the LO-related signals from c1ass directly, as I am not clear why my approach did not work.
More details to follow.

|
8982
|
Wed Aug 7 22:18:43 2013 |
Koji | Update | ASC | ASS update |
While Gautam is working on the Xarm green ASS...
The EPICS monitor points for the ASS actuators were added to the ASS model.
This will be used for the offloading the ASS actuations to the alignment biases.
As this modification allowed us to monitor the actuation apart from the dithering,
now we can migrate the ASS actuation to the fast alignment offset on the suspension.
This modification to the offset moving scripts were also done.
|
8981
|
Wed Aug 7 21:52:11 2013 |
Jenne | Update | SUS | SRM coils fine - problem with slow bias slider |
[Koji, Jenne]
We have looked a little more at the SRM situation. We aligned the SRM, and then aligned the oplev, so that we had a convenient monitor of the optic's motion.
When we use the _COMM channels, which are the usual ones on the IFO_ALIGN screen, the pitch slider makes pitch motion, but the yaw slider makes the oplev spot move ~45degrees from horizontal.
However, when we use the bias channels that are in the front end model, parallel to the ASC path, pitch moves pitch, and yaw moves pure yaw.
So, we conclude that the SRM coils are fine, and there is something funny going on with the slow part of the actuation.
Koji restarted the slow computer susaux, and burt restored it, but that did not fix the situation. We went inside and looked at all of the ribbon cable connections, and pushed them all in, but that also has not fixed things.
We have been looking at D010001-b, the coil driver board, and we think that's where the summing resistor network between the slow bias slider, and the coil outputs from the fast model exists. (It's not 100% clear, but we're confident that that's what is going on).
Tomorrow, we will pull the SRM's coil driver board, and see if any of the components in the slow slider path, before the summing point, look burned / broken / bad. |
8980
|
Wed Aug 7 19:16:20 2013 |
Jenne | Update | ASC | ASS setting up accelerated more |
I have furthered Koji's work, and moved the filter on/off state for all the filter banks also to the burt snapshot.
Turning on the ASS is now much faster than it was originally, with the ezcawrites in series. |
8979
|
Wed Aug 7 15:51:53 2013 |
Alex Cole | Configuration | Electronics | RF Switch Change |
For the photodetector frequency response project, our new RF Switch Chassis (NI pxie-1071) arrived today. I took the switches out of the old chassis (Note for future generations: you have to yank pretty darn hard) and put them in the new chassis, which I mounted in rack 1Y1 as pictured.
The point of this new chassis is that its controller is compatible with our control room computer setup. We will be able to switch the chassis using TCP/IP or telnet, aiding in our automation of the measurement of photodetector frequency response. |
Attachment 1: photo_(2).JPG
|
|
8978
|
Wed Aug 7 15:36:29 2013 |
Steve | Update | VAC | TP2 drypump replaced |
TP2's fore line - dry pump replaced at performance level 600 mTorr after 10,377 hrs of continuous operation.
Where are the foreline pressure gauges? These values are not on the vac.medm screen.
The new tip seal dry pump lowered the small turbo foreline pressure 10x
TP2fl after 2 day of pumping 65mTorr |
Attachment 1: forelinesPs.jpg
|
|
8977
|
Wed Aug 7 15:32:37 2013 |
Koji | Update | ASC | ASS setting up accelerated (slightly) |
I moved bunch of ezcawrite from the ASS Dither On script to a snapshot file.
This accelerated a half of the "up" time but still switching part is not in the snapshot.
If you find anything wrong with ASS, please notify me. |
8976
|
Wed Aug 7 10:51:49 2013 |
Steve | Update | VAC | Pumpdown at 26hrs |
Cold cathode gauge CC1 -h (horizontal) just coming on 9.2e-5 Torr
P2 is the fore line pressure of the maglev. One can see the 4 Torr load during switching over to turbo pumping.
CC4 5e-9 Torr is the performance of the maglev pumping on the RGA only.
The annuloses are not pumped now. They are closed off to see how much outgassing plus leak they have.
Configuration: vacuum normal, annuloses not pumped
Condition: normal
Precondition: 14 days at atm, IOO chamber north door was taken off as a new entrance, the ETMX chamber was not opened.
What is new in the vacuum system: new P1 pirani gauge, gold plated clean allen wrench and ..........what else was dropped?
Note: the wireless laptop did not fail once all day yesterday. I want to give credit to the person who is responsible for this.
|
Attachment 1: 26hr.png
|
|
Attachment 2: 14datm.png
|
|
8975
|
Wed Aug 7 10:09:30 2013 |
Koji | Update | LSC | Arms locked in IR, aligned. IFO at nominal power |
I have a concern about the SRM suspension. The yaw alignment bias produces huge pitch coupling.
This could be a connector issue or the rubbing of the mirror on the EQ stops.
We have the photos of the magnets and they were not touching the OSEMs. |
8974
|
Tue Aug 6 19:53:15 2013 |
Jenne | Update | LSC | Arms locked in IR, aligned. IFO at nominal power |
[Koji, Manasa, Jenne]
The Y arm was locked in IR, and we saw flashing in the Xarm (Gautam had the Xarm for green work when we began). I checked IPANG, and the beam was beautifully unclipped, almost perfectly centered on the first out of vacuum mirror. I aligned the beam onto the QPD.
We then swapped out the MC Y1 that we use at low power, and replace the usual 10% BS, so that we wouldn't crispy-fry MC REFL. Manasa adjusted the half wave plate after the laser, to maximize the power going toward the PMC. We relocked the PMC, and see transmission of ~0.84, which is at the high side of what we usually get. The beam was aligned onto MC REFL and centered on the WFS, and the MC was locked at nominal power. Koji tweaked up the alignment of the MC, and ran the WFS offset script. I aligned beam onto POP QPD and POP110 coarsely (using a flashing PRC, not a locked PRM-ITMY cavity, so the alignment should be rechecked). The arms have both been locked and aligned in IR....the green beams need to be steered to match the current cavity axis.
The AS beam, as well as REFL and POP, are all coming out of the vacuum nicely unclipped.
Notes: When Koji was aligning the SRM to get the SRC cavity roughly aligned (the AS flashes all overlapping), we noticed that there is some major pitch-yaw coupling. Serious enough that we should be concerned that perhaps some connector is loose, or an actuator isn't working properly. This should be checked.
Moral of the story: Coarse alignment of all mirrors is complete after pump-down and we have IR locked and aligned to both arms at nominal power.
Still to do:
* Restore PRM, align beam onto the REFL PDs.
* Lock PRM-ITMY cavity, align beam onto POP PDs.
* Align AS beam onto AS55.
* Recenter all oplevs.
* Recenter IPPOS and IPPANG at nominal power.
* Start locking!! |
8973
|
Tue Aug 6 16:45:31 2013 |
Steve | Update | VAC | Pumpdown is completed |
IFO P1=1mT PSL shutter is opened
|
Attachment 1: 8hrPd76.png
|
|
Attachment 2: 1mT.png
|
|
8972
|
Tue Aug 6 16:36:51 2013 |
gautam | Update | CDS | Choosing LO Amplitudes and Frequencies-revised |
I redid the power spectrum measurement for the X-arm green transmission after aligning the arm to green using the ITMX/ETMX Pitch and Yaw sliders on IFOalign.
The Y-axis now reflects the relative intensity noise (RIN), which I obtained by taking the average value of the X-arm green transmission using tdsavg. Based on this measurement, I have now picked four new frequencies at which to try and modulate the PZT mirrors: 10, 19, 34 and 39 Hz. Bandpass filters in the LIA stage have been appropriately modified.
Power Spectrum:

|
8971
|
Tue Aug 6 12:43:23 2013 |
Alex Cole | Configuration | Electronics | AS Table and Rack 1Y1 Additions |
For the photodetector frequency response project, I finished the construction of our baluns chassis and mounted it in rack 1Y1 (1st picture).
After consulting with Jenne, I mounted the fiber launcher for REFL165 on the AS table such that it would not cause an obstruction. I aligned the launcher using a multimeter to monitor the DC output of REFL165, but looking at the data I got, it seems I need to do a better alignment/focusing job to get rid of a bunch of noise. |
Attachment 1: photo_1_(5).JPG
|
|
Attachment 2: photo_2_(5).JPG
|
|
8970
|
Tue Aug 6 10:06:29 2013 |
Steve | Update | VAC | Pumpdown is at 500Torr |
The pumpdown has started at 8:38am
Manasa was here to confirm good alignment of the IFO
Inner jam nuts of AC bellow were torqued to 45 ft/lbs and door nuts were check on opened chambers.
Annulos were roughed down to 500 mTorr
Oplev servos turned off, PSL and green shutters closed before pumpdown started
|
Attachment 1: befPd.png
|
|
8969
|
Tue Aug 6 07:52:58 2013 |
manasa | Update | General | Ready to pump down |
I did an alignment check of the IFO before we start pumping down.
Arms were locked. PRM and SRM were aligned. Green was aligned to the arms for reference during the pump down.
Steve! It's a GO!
MC spot positions:

OSEM:

Oplevs were all centered yesterday and haven't drifted much. So I left them as is.

QPDs (IPPOS aligned from yesterday.

|